OpenStudy (anonymous):

kingkos Group Neophyte function help given f(x) = 4√-x find: 1. f(4-x) 2. f(4-x^2) 3. f(4-x^4) 4. f(4x- x^2) 5. f(-x^4 - 4x^2) pls help me for my H.w

OpenStudy (anonymous):

replace x by the inputs and make sure to use parentheses

OpenStudy (anonymous):

\[f(x)=\sqrt[4]{-x}\]

OpenStudy (anonymous):

thats the given sir

OpenStudy (anonymous):

\[f(4-x)=\sqrt[4]{-(4-x)}=\sqrt[4]{(x-4)}\]

OpenStudy (anonymous):

thats for no.1 sir? ic i just connect or simplify..

OpenStudy (anonymous):

\[f(4-x^2)=\sqrt[4]{-(4-x^2)}=\sqrt[4]{x^2-4}\]\]

OpenStudy (anonymous):

there is nothing to "simplify" here. just write what you get

OpenStudy (anonymous):

great yes!

OpenStudy (anonymous):

3 more to go yehey!

OpenStudy (anonymous):

so next one is \[f(4-x^4)=\sqrt[4]{x^4-4}\]

OpenStudy (anonymous):

then \[f(4x-x^2)=\sqrt[4]{x^2-4x}\]

OpenStudy (anonymous):

and finally \[f(-x^4-4x^2)=\sqrt[4]{x^4+4x^2}\]

OpenStudy (anonymous):

god bless u sir thank thank you!